# Fully stochastic trust-region methods with Barzilai-Borwein steplengths

### v1

@@ 2.3 TRishBB_v3
Algorithm TRishBB_v3 is originated by the arguments in [32] to avoid both the potentially harmful effects of the gradient difference when $\|s_{k}\|$ is small and the noise in stochastic gradient estimates. In [32], given $s_{k}=\bar{x}-\bar{x}_{old}$ for some $\bar{x},\,\bar{x}_{old}$, the vector $y_{k}$ is defined as $y_{k}=\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})s_{k}$ where $\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})$ is a subsampled Hessian approximation and $|\widetilde{\mathcal{N}_{k}}|$ is supposed to be larger than the batch size used for gradients. If the objective function $f$ is the cross-entropy loss function or the least-squares function, then $\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})$ can be approximated by the empirical Fisher Information Matrix (FIM) [33, §11] defined as
$\frac{1}{|\mathcal{\bar{N}}_{k}|}\sum_{i\in\widetilde{\mathcal{N}_{k}}}\nabla f_{i}(\bar{x})\nabla f_{i}(\bar{x})^{T}.$
Empirical FIM has a higher per-iteration cost than first-order methods, thus it was further simplified by introducing the accumulated empirical FIM [34, 35].
In Algorithm TRishBB_v3, we update $s_{k}$ and $y_{k}$ every $m$ iterations and $y_{k}$ is computed as the accumulated empirical FIM. In particular, as displayed in line 13 of the algorithm, $s_{k}$ is formed using the averages $\bar{x}$ and $\bar{x}_{old}$; $\bar{x}$ is the average of the last $m$ iterates, and $\bar{x}_{old}$ is the average of the previous $m$ iterates, see lines 2 and 11 of the algorithm. Average iterates are meant to represent more stable approximations than a single iterate for computing the difference $s_{k}$. Concerning $y_{k}$, let $F_{k}$ be the limited memory matrix whose columns contain the last $L$ computed stochastic gradients for some positive integer $L$. Then, the curvature vector $y_{k}$ is defined as ${y}_{k}=\frac{1}{|F_{k}|}F_{k}(F_{k}^{T}s_{k})$, see line 13 of the algorithm, where $|F_{k}|$ denotes the number of columns of $F_{k}$. Since $F_{k}F_{k}^{T}$ is the sum of $|F_{k}|$ rank-one matrices, (6) implies
$\mu_{k+1}^{\text{BB}}=\frac{1}{\frac{1}{|F_{k}|}\sum_{t=k-|F_{k}|+1}^{k}q(\nabla f_{\mathcal{N}_{t}}(x_{t})(\nabla f_{\mathcal{N}_{t}}(x_{t}))^{T},{s}_{k})},$
where $\nabla f_{\mathcal{N}_{t}}(x_{t})(\nabla f_{\mathcal{N}_{t}}(x_{t}))^{T}$ is a rank-one matrix.
@@ 3.2 Analysis with Unbiased Gradients
Hence Assumption (3.3) is satisfied with $\omega=1$, $M_{1}=M_{g}$, $M_{2}=1$ and the convergence results of §3.1 hold. Now we further characterize the convergence without imposing bounds, such as inequality (18), on $\gamma_{1}$, $\gamma_{2}$. The first lemma is a technical result.
@@
When $f$ is a nonconvex funtion, the expected average squared norm of the gradient converges to $\left(\frac{16\gamma_{1}^{2}}{\gamma_{2}^{2}}-1\right)M_{g}$.
@@ 4 Numerical Experiments
In this section, we present experimental results to compare Algorithms TRishBB_v1 , TRishBB_v2 and TRishBB_v3 with the first-order version of TRish in [7], i.e., Algorithm TRishBB with $\mu_{k}=0$, $\forall k$. All experiments were conducted with MATLAB (R2024a) on a Rocky Linux 8.10 (64-bit) server with 256 GiB memory using an NVIDIA A100 PCIe GPU (with 80 GiB on-board memory).
@@ 4.1 Classification Problems
Table 1: Classification test problems
Dataset $\mathbf{(N,\bar{N})}$ $\mathbf{d}$ $\mathbf{n}$ Type $\mathbf{f_{i}(x)}$ NN
a1a [36] $(1605,29351)$ $123$ $123$ Binary Logistic regression -
w1a [36] $(2477,47272)$ $300$ $300$ Binary Logistic regression -
cina [36] $(10000,6033)$ $132$ $132$ Binary Logistic regression -
MNIST [37] $(60000,10000)$ $28\times 28\times 1$ $431030$ Multi Cross-entropy NET-1
CIFAR10 [38] $(50000,10000)$ $32\times 32\times 3$ $545098$ Multi Cross-entropy NET-2
We consider some ML test problems of the form (1). Table 1 summarizes five test problems used in our experiments for both binary and multi-class classification tasks. Each test dataset with $C$ possible classes is divided into training and testing sets, denoted as $\{(a_{i},b_{i})\,|i\in\mathcal{N}\}$ and $\{(a_{i},b_{i})\,|i\in\mathcal{\bar{N}}\}$, with $|\mathcal{N}|=N$ and $|\mathcal{\bar{N}}|=\bar{N}$ samples, respectively; the values of $N$ and $\bar{N}$ are reported in the table. The vector $a_{i}\in\mathbb{R}^{d}$ represents the $i$-th data with $d$ features while $b_{i}$ is the corresponding categorical label associated. The dataset cina was normalized such that each feature vector has components in the interval $[0,1]$. For MNIST and CIFAR10, besides the zero-one rescaling, we implemented z-score normalization to have zero mean and unit variance. Precisely, all $N$ training images, (3-D arrays of size $h\times w\times c$ where $h$, $w$, and $c$ denote the height, width, and number of channels of the images, respectively), undergo normalization by subtracting the mean (a 3-D array) and dividing by the standard deviation (a 3-D array) component-wise. Data in the testing set are also normalized using the same statistics as in the training data.
For binary classification problems, we used logistic regression where the single loss function $f_{i}(x)$ with $x\in\mathbb{R}^{n}$ in (1) is defined as
@@
This model is used for the datasets a1a, w1a, and cina. Note that the number of unknown parameters in $x\in\mathbb{R}^{n}$ is equal to the number of features in $a_{i}\in\mathbb{R}^{d}$, i.e., $n=d$. For multi-classification problems, we used neural network models coupled with the softmax cross-entropy loss that is the single loss function $f_{i}(x)$ with $x\in\mathbb{R}^{n}$ in (1), and defined as
$f_{i}(x)=-\sum_{k=1}^{C}(b_{i})_{k}\log(h(a_{i};x))_{k},\quad i=1,\dots,N.$
We employ this model using the datasets MNIST and CIFAR10 where $a_{i}\in\mathbb{R}^{d}$ is a 3-D array (grayscale or RGB image), and the categorical label $b_{i}$ is converted into a one-hot encoded vector; i.e., it is a binary vector of length $C$, where $C-1$ elements are zero except for the element whose index corresponds to the true class that is equal to one. The output $h(a_{i};\cdot)$ is a prediction provided by the neural network (NN) whose architecture is described in Table 2.111The compound $Conv(5,20,1,0){/}ReLU{/}MaxPool(2,2,0)$ indicates a convolutional layer ($Conv$) using $20$ filters of size $5\times 5$, stride $1$, and no padding, a nonlinear activation function ($ReLu$) and, a 2-D max-pooling layer with a channel of size $2\times 2$, stride 2 and no padding. The syntax $FC(C){/}Softmax$ denotes a fully connected layer of $C$ neurons followed by the $softmax$ layer. For MNIST dataset with grayscaled images in $C=10$ classes (digits 0–9), and CIFAR10 with RGB images in $C=10$ categories (e.g., dogs and ships), we used NET-1 and NET-2 respectively. Note that the size $n$ of problem (1) using an NN depends on both the number of features $d$ and the depth of the network. We utilized the MATLAB Deep Learning Toolbox, leveraging automatic gradient computation in customized training loops; we used forward and crossentopy for implementing the NN softmax cross-entropy. See [39, Table2] for details on implementing NN classification problems by this toolbox.
Table 2: NN architecture
NET-1 $Input(d)$
$Conv(5,20,1,0){/}ReLU{/}MaxPool(2\times 2,2,0)$
$Conv(5,50,1,0){/}ReLU{/}MaxPool(2\times 2,2,0)$
$FC(500){/}ReLU$
$FC(C){/}Softmax$
NET-2 $Input(d)$
$Conv(3,32,1,same){/}ReLU{/}AvragePool(2\times 2,2,0)$
$Conv(3,64,1,same){/}ReLU{/}AvragePool(2\times 2,2,0)$
$FC(128){/}ReLU$
$FC(C){/}Softmax$
@@ 4.2 Experimental Configuration
Algorithm TRishBB_v1 was run setting $m=20$ and selecting random and independent mini-batches ${\cal{N}}_{k}$, thus the gradients $g_{k}$ are unbiased. Algorithm TRishBB_v2 was run with $\eta=0.9$ and $m=N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ on the base of the discussion in §2.2; the mini-batches ${\cal{N}}_{k}$ were selected as disjoint subsets of ${\cal{N}}$ and the resulting gradients are biased. Algorithm TRishBB_v3 was run setting $m=5$, $L=100$ as suggested in [34] and the mini-bath selection was as in TRishBB_v1. For all $k\geq 0$, the mini-batch size for computing stochastic gradients in these variants was $|\mathcal{N}_{k}|=64$ for the datasets a1a, w1a, and cina and $|\mathcal{N}_{k}|=128$ for the datasets MNIST and CIFAR10. The values $\mu_{\min}$, and $\mu_{\max}$ were set to $10^{-5}$ and $10^{5}$, respectively.
Our experiments were carried out as in [7]. We ran the algorithms with varying settings of triplets $(\alpha,\gamma_{1},\gamma_{2})$ whose values are given in Table 3 for each dataset. In the table we also report the value of $G$ used for defining the triplets, that is the average norm of the stochastic gradients generated in a one-epoch run of the SGD method with a fixed learning rate $\ell$. We used 60 and 36 triplets for binary classification and multi-class problems, respectively. For the binary and multi-class test problems, given the triplet $(\alpha,\gamma_{1},\gamma_{2})$, we ran each algorithm 50 and 10 times using 50 and 10 predetermined random seeds, respectively. We present the testing accuracy, i.e., the percentage of samples in the testing set that the model correctly classifies; in case of multiple runs, the testing accuracy is averaged over the runs. In the figures shown in the following subsection, we use ExIJK to denote the results obtained in the Experiment where $\alpha$, $\gamma_{1}$, and $\gamma_{2}$ take the I-th, J-th, and K-th value indicated in Table 3, respectively.
In the solution of problems with logistic regression objective function, the initial guess was the zero vector, $x_{0}=0$. For solving the NN softmax cross-entropy problems, we used the Glorot (Xavier) initializer [40] for the weights, and zeros for the biases in the network’s layers. All algorithms terminated after five number of epochs, namely when the total number of single gradient evaluations $\nabla f_{i}$, reached or exceeded $5N$, where $N$ is the number of training samples.
Table 3: Hyper-parameters for the experiments
@@ 4.3 Numerical Results
In Table 5, averaging over the tested values of $\gamma_{1}$ and $\gamma_{2}$, for every dataset and every value of $\alpha$ we report the highest accuracy obtained by TRish and the variants of TRishBB; the reported values represent the largest accuracies among those reached at the end of each epoch. For each algorithm we also report the ratio (“accuracy ratio”) between the smallest and the largest accuracy reported in the corresponding line of the table. We can observe that the accuracy of TRishBB variants is comparable or higher than the accuracy of TRish and that the best results for TRishBB are obtained with the largest values of $\alpha$ where it is possible to benefit of the BB steplenghts for a large number of iterations. In order to give more insight on the behaviour of our three TRish variants in what follows we provide detailed statistics of the runs employing the largest value of $\alpha$.
Table 4: Percentage of BB steplengths taken with different $\alpha$
@@
TRishBB_v3 00.00 00.00 00.00 01.84
Table 5: Highest testing accuracy with different $\alpha$ (in %) and the accuracy ratio
@@ 4.3.1 Datasets a1a, w1a, and cina
We present the averaged accuracies obtained with $\alpha=10$ for each of the five epochs performed in the experiments and for each triplet $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$.
@@
Figure 1Figure 3 illustrate the results obtained on the dataset a1a. The results displayed indicate that the highest accuracy is achieved by TRishBB_v2 in the first two epochs. The algorithms TRishBB_v1 and TRishBB_v2 outperform TRish, remarkably the testing accuracy reached is quite insensitive to the tested values of $\gamma_{1}$ and $\gamma_{2}$. For large $\alpha$, these algorithms take the unconstrained step in the large majority of the iterations, see Table 4, and thus differ from TRish considerably. On the other hand, TRishBB_v3 is generally more accurate than TRish but more sensitive to the choice of $\gamma_{1}$ and $\gamma_{2}$; the steplength $\mu_{k}$ was selected approximately in $60\%$ of the iterations.
@@
Figure 1: a1a: Average accuracy of TRishBB_v1 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=20$, and $|\mathcal{N}_{k}|=64$.
Figure 2: a1a: Average accuracy of TRishBB_v2 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=N_{b}$, $|\mathcal{N}_{k}|=64$.
Figure 3: a1a: Average accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=5$, and $|\mathcal{N}_{k}|=64$.
Figure 4Figure 6 concern the dataset w1a. Algorithm TRishBB_v2 was very effective and its performance is insensitive to the triplet $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$ used. Using this triplet, the steplength $\mu_{k}$ was selected at every iteration by this algorithm, see Table 4. On the other hand, the steplength $\mu_{k}$ was selected approximately $79\%$ and $16\%$ of the iterations by TRishBB_v1 and TRishBB_v3respectively. Algorithms TRishBB_v1 and TRishBB_v3 provided classification accuracies comparable to those of TRish.
Results for the dataset cina are displayed in Figure 7Figure 9. From Table 4 we know that TRishBB_v1 selected the steplength $\mu_{k}$ approximately at the $89\%$ of the iterations, TRishBB_v2 selected $\mu_{k}$ at all iterations while TRishBB_v3 selected $\mu_{k}$ around at the $72\%$ of all iterations. TRishBB_v1 is generally more accurate than TRish but more sensitive to the choice of $\gamma_{1}$ and $\gamma_{2}$ than TRishBB_v2; Algorithm TRishBB_v2 achieves high accuracy steadily for all the values $\gamma_{1}$ and $\gamma_{2}$. Algorithm TRishBB_v3 is the most sensitive to the values of $\gamma_{1}$ and $\gamma_{2}$. From the second to the fifth epoch, it compares well with TRish but the accuracy is generally lower than the accuracy reached by TRishBB_v2.
@@
The results described above indicate that exploiting stochastic BB steplengths and guidelines on stochastic quasi-Newton methods seems to considerably improve the overall performance of TRish. Algorithm TRishBB_v2 provides the best accuracy and its performance is not affected by the specific triplet used. Figure 10 shows the average values of the BB parameters built with varying $\gamma_{1}$ and $\gamma_{2}$ for each dataset. It is evident that the rule for generating $\mu_{k}$ in TRishBB_v2 provides small values, thus enhancing the selection of unconstrained steps $(-\mu_{k}g_{k})$. Coherently with the results above, the largest $\mu_{k}$’s are built in TRishBB_v3.
@@
Figure 4: w1a: Average accuracy of TRishBB_v1 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=20$, $|\mathcal{N}_{k}|=64$.
Figure 5: w1a: Average accuracy of TRishBB_v2 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=N_{b}$, $|\mathcal{N}_{k}|=64$.
Figure 6: w1a: Average accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=5$, $|\mathcal{N}_{k}|=64$.
@@
Figure 7: cina: Average accuracy of TRishBB_v1 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=20$, $|\mathcal{N}_{k}|=64$.
Figure 8: cina: Average accuracy of TRishBB_v2 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=N_{b}$, $|\mathcal{N}_{k}|=64$.
Figure 9: cina: Average accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m=5$, $|\mathcal{N}_{k}|=64$
@@
Figure 10: Average values of $\mu_{k}$ with triplets $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$ along iterations for a1a, w1a and cina.

### v2

@@ 2.3 TRishBB_v3
Algorithm TRishBB_v3 is originated by the arguments in [32] to avoid both the potentially harmful effects of the gradient difference when $\|s_{k}\|$ is small and the noise in stochastic gradient estimates. In [32], given $s_{k}=\bar{x}-\bar{x}_{old}$ for some $\bar{x},\,\bar{x}_{old}$, the vector $y_{k}$ is defined as $y_{k}=\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})s_{k}$ where $\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})$ is a subsampled Hessian approximation and $|\widetilde{\mathcal{N}_{k}}|$ is supposed to be larger than the batch size used for gradients. If the objective function $f$ is the cross-entropy loss function or the least-squares function, then $\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})$ can be approximated by the empirical Fisher Information Matrix (eFIM) [33, §11] defined as
$\frac{1}{|\widetilde{\mathcal{N}_{k}}|}\sum_{i\in\widetilde{\mathcal{N}_{k}}}\nabla f_{i}(\bar{x})\nabla f_{i}(\bar{x})^{T}.$
In order to reduce the cost, the eFIM was further simplified by introducing the accumulated Fisher Information Matrix (aFIM) [34, 35]. In Algorithm TRishBB_v3, we update $s_{k}$ and $y_{k}$ every $m$ iterations, and $y_{k}$ is computed using the aFIM and the first-in, first-out (FIFO) approach [34, 35]. In particular, as displayed in line 13 of the algorithm, $s_{k}$ is formed using the averages $\bar{x}$ and $\bar{x}_{old}$; $\bar{x}$ is the average of the last $m$ iterates, and $\bar{x}_{old}$ is the average of the previous $m$ iterates, see lines 2 and 11 of the algorithm. Average iterates are meant to represent more stable approximations than a single iterate for computing the difference $s_{k}$. Concerning $y_{k}$, let $F_{k}$ be the limited memory matrix whose columns contain at most the last $m_{F}$ computed stochastic gradients for some positive integer $m_{F}$. Then, the curvature vector $y_{k}$ is defined as ${y}_{k}=\frac{1}{|F_{k}|}F_{k}(F_{k}^{T}s_{k})$, see line 13 of the algorithm, where $|F_{k}|$ denotes the number of columns of $F_{k}$. Since $F_{k}F_{k}^{T}$ is the sum of $|F_{k}|$ rank-one matrices, (6) implies
$\widehat{\mu}_{k}=\frac{|F_{k}|}{m}\frac{1}{\sum_{t=k-|F_{k}|+1}^{k}q(\nabla f_{\mathcal{N}_{t}}(x_{t})(\nabla f_{\mathcal{N}_{t}}(x_{t}))^{T},{s}_{k})},$
where $\nabla f_{\mathcal{N}_{t}}(x_{t})(\nabla f_{\mathcal{N}_{t}}(x_{t}))^{T}$ is a rank-one matrix. Finally, the update of $\mu_{k}$ is performed every $m$ iterations and it is the same as in TRishBB_v2.
Compared to TRish, the extra cost of this algorithm amounts to $m_{F}+2$ scalar products to evaluate $\mu_{k+1}$; in terms of storage, two vectors of size $n$ and one matrix of size $n\times m_{F}$ are required.
@@ 3.2 Analysis with Unbiased Gradients
Hence Assumption (3.3) is satisfied with $\omega=1$, $M_{1}=M_{g}$, $M_{2}=1$ and the convergence results of Section 3.1 hold. Now we further characterize the convergence without imposing bounds, such as inequality (18), on $\gamma_{1}$, $\gamma_{2}$. The first lemma is a technical result.
@@
When $f$ is a nonconvex function, the expected average squared norm of the gradient converges to $\left(\frac{16\gamma_{1}^{2}}{\gamma_{2}^{2}}-1\right)M_{g}$.
@@ 4 Numerical Experiments
In this section, we present experimental results to compare Algorithms TRishBB_v1, TRishBB_v2 and TRishBB_v3 with the first-order version of TRish in [7], i.e., Algorithm TRishBB with $H_{k}=0$, $\forall k$. All experiments were conducted with MATLAB (R2024a) on a Rocky Linux 8.10 (64-bit) server with 256 GiB memory using an NVIDIA A100 PCIe GPU (with 80 GiB on-board memory).
@@ 4.1 Classification Problems
We consider some ML and DL test problems of the form (1). For binary classification problems using a1a, w1a, and cina datasets [36], we used logistic regression where the single loss function $f_{i}(x)$ with $x\in\mathbb{R}^{n}$ in (1) is defined as
@@
Note that $x\in\mathbb{R}^{n}$ and the data sample $a_{i}\in\mathbb{R}^{d}$ with the label $b_{i}$ have the same dimension, i.e., $n=d$. A numerically stable implementation of the logistic sigmoid was employed to prevent numerical overflow.
For multi-classification problems using MNIST [37] and CIFAR10 [38] datasets with $C$ classes, we used neural network models coupled with the softmax cross-entropy loss in (1), whose single loss function $f_{i}(x)$ is defined as
$f_{i}(x)=-\sum_{k=1}^{C}(b_{i})_{k}\log(h(a_{i};x))_{k},\quad i=1,\dots,N,$
where $a_{i}\in\mathbb{R}^{d}$ is a 3-D array (grayscale or RGB image), and the categorical label $b_{i}$ is converted into a one-hot encoded vector; i.e., it is a binary vector of length $C$, where $C-1$ elements are zero except for the element whose index corresponds to the true class that is equal to one. The output $h(a_{i};\cdot)$ is a prediction provided by the neural network (NN) whose architecture is described in Table 8 in Appendix C. The size $n$ of problem (1) using a NN depends on both the number of features $d$ and the depth of the network. We utilized the MATLAB Deep Learning Toolbox, leveraging automatic gradient computation in customized training loops; we used forward and crossentropy for implementing the NN softmax cross-entropy. See [39, Table2] for details on implementing NN classification problems by this toolbox.
More details on the test datasets and the architecture of the NNs used are available in Appendix C.
@@ 4.2 Experimental Configuration
Algorithm TRishBB_v1 was run setting $m=20$ and selecting random and independent mini-batches ${\cal{N}}_{k}$, thus the gradients $g_{k}$ are unbiased. Algorithm TRishBB_v2 was run with $\eta=0.9$ and $m=N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ on the base of the discussion in Section 2.2; the mini-batches ${\cal{N}}_{k}$ were selected as disjoint subsets of ${\cal{N}}$, with ${\cal{N}}$ being shuffled at the beginning of each epoch; thus, the resulting gradients are biased. Algorithm TRishBB_v3 was run with $m_{F}=100$ as suggested in [34], its mini-batch selection was as in TRishBB_v1, and its cyclic iteration parameter $m$ was selected as in TRishBB_v2.
For all $k\geq 0$, the mini-batch size for computing stochastic gradients in these variants was $|\mathcal{N}_{k}|=64$ for the datasets a1a, w1a, and cina and $|\mathcal{N}_{k}|=128$ for the datasets MNIST and CIFAR10. The initial BB steplength $\mu_{0}$ was set to 1 in all the algorithms; moreover, the values $\mu_{\min}$, and $\mu_{\max}$ were set to $10^{-5}$ and $10^{5}$, respectively.
Regarding the predefined parameters $(\alpha,\gamma_{1},\gamma_{2})$, our settings were borrowed from those in [7] which are accompanied by a detailed discussion. The scalar $\alpha$ takes value in a prefixed set and the triplets $(\alpha,\gamma_{1},\gamma_{2})$ used in our experiments are given in Table 1 for each dataset. The parameters $\gamma_{1}$ and $\gamma_{2}$ are chosen by a gradient estimation as follows. We run the SG algorithm for one epoch with a fixed learning rate $\ell$ and compute the norm of the stochastic gradient at each iteration; then we set $G$ as the average of the norms computed throughout the epoch. The value $G$ estimates the magnitude of the stochastic gradients and promotes the occurrence of the three cases in (8) and (9), i.e., the use of normalized steps. Overall we used 60 and 36 triplets for binary classification and multi-class problems, respectively.
For the binary and multi-class test problems, given the triplet $(\alpha,\gamma_{1},\gamma_{2})$, we ran each algorithm 50 and 10 times using 50 and 10 predetermined random seeds, respectively. We present the testing accuracy, i.e., the percentage of samples in the testing set correctly identified; in case of multiple runs, the testing accuracy is averaged over the runs. In the figures shown in the following subsection, we use ExIJK to denote the results obtained in the Experiment where $\alpha$, $\gamma_{1}$, and $\gamma_{2}$ take the I-th, J-th, and K-th value indicated in Table 1, respectively.
[…]
Table 1: Hyper-parameters for the experiments.
@@ 4.3 Numerical Results
In Table 3, we report, for every dataset and every value of $\alpha$, the highest testing accuracy obtained by TRish and the variants of TRishBB averaging over the tested values of $\gamma_{1}$ and $\gamma_{2}$; the reported values represent the largest testing accuracies among those reached at the end of each epoch. For each algorithm we also report the ratio (”accuracy ratio”) between the smallest and the largest accuracy reported in the corresponding line of the table. We can observe that the highest testing accuracy of the TRishBB variants is comparable to or larger than the accuracy of TRish. For the datasets a1a, w1a, and cina, the values of the testing accuracy ratio are close to one for TRishBB_v2 and TRishBB_v3, namely the highest testing accuracy obtained over five epochs is not highly sensitive to the value of $\alpha$. In contrast, the testing accuracy reached by the three versions of TRishBB and TRish on the DL test problems is dependent on $\alpha$. In order to give more insight, in what follows we provide detailed statistics of the runs employing the largest value of $\alpha$ which promotes the use of the BB steplengths.
Table 2: Percentage of BB steplengths taken with different $\alpha$.
@@
TRishBB_v3 00.00 00.00 00.00 00.78
Table 3: Highest average testing accuracy with different $\alpha$ (in %) and the accuracy ratio.
@@ 4.3.1 Datasets a1a, w1a, and cina
We present the averaged accuracies obtained with $\alpha=10$ for each of the five epochs and for each triplet $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$. For this value of $\alpha$, our algorithms take the unconstrained step in the large majority of the iterations, and thus differ from TRish considerably, see Table 2.
@@
Figure 1 illustrates the results obtained on the dataset a1a. The results displayed indicate that the highest accuracy is achieved by TRishBB_v2 and TRishBB_v3 in the first two epochs and that all TRishBB variants outperform TRish. Remarkably the testing accuracy reached by TRishBB is quite insensitive to the tested values of $\gamma_{1}$ and $\gamma_{2}$ contrary to TRish.
@@
Figure 1: a1a: Average accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=64$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=25$. Bottom: TRishBB_v3, $m=25$.
[…]
@@
The results described above indicate that exploiting stochastic BB steplengths and guidelines on stochastic quasi-Newton methods improves the overall performance of TRish. Figure 4 shows the average values of the BB parameters built with $\alpha=10$ and varying $\gamma_{1}$ and $\gamma_{2}$ for each dataset. It is evident that the rule for generating $\mu_{k}$ in TRishBB_v2 and TRishBB_v3 provides smaller values of it, enhancing the selection of unconstrained steps $(-\mu_{k}g_{k})$. Algorithm TRishBB_v2 provides the best accuracy and its performance is quite insensitive to the triplet used.
@@
Figure 2: w1a: Average accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=64$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=38$. Bottom: TRishBB_v3, $m=38$.
@@
Figure 3: cina: Average accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=64$. Top: TRishBB_v1, $m=20$; Middle: TRishBB_v2, $m=156$. Bottom: TRishBB_v3, $m=156$.
@@
Figure 4: Average values of $\mu_{k}$ with triplets $(\alpha,\gamma_{1},\gamma_{2})$ along the iterations. Top: a1a, w1a and cina with $\alpha=10$. Bottom: MNIST and CIFAR10 with $\alpha=1$.